OpenStudy (anonymous):

Prove (cscx+cotx)(cscx-cotx)=1 ((1/sinx)+(cos/sinx))(1/sinx-(cosx/sinx)) (1+cosx/sinx)(1-cosx/sinx) I don't understand what to do next or if I even Have it right

OpenStudy (accessdenied):

my immediate reaction from the beginning would have been "difference of squares! csc^2 x - cot^2 x = 1" i cant remember any identities about those functions tho

OpenStudy (accessdenied):

i guess if you use 1= sin^2 x / sin^2 x, you can multiply everything by sin^2 x and get 1 - cos^2 x = sin^2 x which can be manipulated to a known identity

jimthompson5910 (jim_thompson5910):

You can use the idea that cos^2(x) + sin^2(x) = 1. Then divide all terms by sin^2(x) to get cos^2(x)/sin^2(x) + sin^2(x)/sin^2(x) = 1/sin^2(x) cot^2(x) + 1 = csc^2(x) 1 = csc^2(x) - cot^2(x) csc^2(x) - cot^2(x) = 1 So csc^2(x) - cot^2(x) = 1 is an identity

OpenStudy (anonymous):

Thank You :))

jimthompson5910 (jim_thompson5910):

you're welcome